OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? x^4-41x^2=-400

OpenStudy (anonymous):

oh no i made a mistake, it is not that bad at all!!

OpenStudy (anonymous):

\[x^4-41x^2+400=0\] \[(x^2-16)(x^2-25)=0\] \[(x-4)(x+4)(x-5)(x+5)=0\] solve for \(x\)